# Cohomology of Lie algebroids over Algebraic spaces

### v6

@@ 2.2. Universal enveloping algebroid of a Lie algebroid
Let $(\mathcal{L},[\cdot,\cdot],\mathfrak{a})$ be a Lie algebroid over any of the special $a$-space $(X,\mathcal{O}_{X})$. For each open set $U$ of $X$, we find the universal enveloping algebra $\mathcal{U}(\mathcal{O}_{X}(U),\mathcal{L}(U))$ of the $(\mathbb{K},\mathcal{O}_{X}(U))$-Lie-Rinehart algebra $\mathcal{L}(U)$ or of the Lie-Rinehart pair $(\mathcal{O}_{X}(U),\mathcal{L}(U))$ (see [Rin63, Hue90, MM10]). The sheafification of the cannonical presheaf:
@@
is known as the universal enveloping algebroid of the Lie algebroid $\mathcal{L}$, and denoted by $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ (see [CRvdB10, Pym13, Bru17, MS22] for details).
@@
The sheaf of logarithmic derivations and sheaf of logarithmic differential operators are denoted as $\mathcal{T}_{X}(-logY)$ and $\mathcal{D}_{X}(-logY)$ respectively for a principal divisor $Y$ in some complex manifold or smooth algebraic variety $X$ [MS22].
In the case of a free divisor $Y$ in $X$ (i.e. $\mathcal{T}_{X}(-logY)$ is locally free $\mathcal{O}_{X}$-module [CJNMM96, Pym13, MS22]) we have (sheafifying the local description given for the module of logarithmic derivations in [Mac14]) the isomorphism
@@
In [Kap07], the notion of free Lie algebroid $\mathcal{P}_{X}$ (called the path algebroid of a smooth manifold or smooth algebraic variety $X$) is constructed for tangent sheaf $\mathcal{T}_{X}$. It forms a locally free sheaf of Lie-Rinehart algebras over $(X,\mathcal{O}_{X})$ of infinite rank. The universal enveloping algebroid $\mathscr{U}(\mathcal{O}_{X},\mathcal{P}_{X})=:\mathbb{D}_{X}$ is described as sheaf of non commutative differential operators on $X$.
@@ 2.3. $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebras
The notation $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebras [MS22] is considered as a sheaf theoretic analogue of the notion $R/{\mathbb{K}}$-bialgebras [MM10] or left bialgebroids [KP11]. As an example, we consider the universal enveloping algebroid $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ (and its dual) of a Lie algebroid $\mathcal{L}$ over $(X,\mathcal{O}_{X})$. Similar structure is appeared in [Kap07, CVdB10, BP15], in complex and algebraic geometry context.
@@
We recall the canonical $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra structures on $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ and its dual.
@@
Observe that $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ is a cocommutative counital $\mathcal{O}_{X}$-coalgebra in the sense that there is a natural cocommutative coassociative comultiplication
$\Delta:\mathscr{U}(\mathcal{O}_{X},\mathcal{L})\rightarrow\mathscr{U}(\mathcal{O}_{X},\mathcal{L})\otimes_{\mathcal{O}_{X}}\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$, and counit $\epsilon:\mathscr{U}(\mathcal{O}_{X},\mathcal{L})\rightarrow\mathcal{O}_{X}$ respectively, which are locally given by the following formulas (using the Sweedler convention) [CRvdB10]
@@
$\epsilon(D)=D(1)$,
for a section $f$ of $\mathcal{O}_{X}$, for a section $D$ of $\mathcal{L}$ and for sections $D^{\prime},D^{\prime\prime}$ of $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$. Consequently, $(\mathscr{U}(\mathcal{O}_{X},\mathcal{L}),\Delta,\epsilon)$ is a sheaf of $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebras.
@@
For each open set $U$ of $X$, we find the jet algebras $\mathcal{J}(\mathcal{O}_{X}(U),\mathcal{L}(U))$ of the $(\mathbb{K},\mathcal{O}_{X}(U))$-Lie-Rinehart algebra $\mathcal{L}(U)$ (see [BP15, CVdB10, KP11]). The jet algebroid $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$ is the sheafification of the canonical presheaf:
@@
Observe now that the product on $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ descends to a sheaf homomorphism $m:\mathscr{U}(\mathcal{O}_{X},\mathcal{L})\otimes_{\mathcal{O}_{X}}\mathscr{U}(\mathcal{O}_{X},\mathcal{L})\rightarrow\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$. We can therefore dualize the product to obtain a coproduct $\Delta^{*}:\mathscr{J}(\mathcal{O}_{X},\mathcal{L})\rightarrow\mathscr{J}(\mathcal{O}_{X},\mathcal{L})\otimes_{\mathcal{O}_{X}}\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$ locally defined as
@@
The jet algebroid is formal groupoid which serve as the formal exponentiation of the Lie algebroid [CRvdB10]. In [Kap07], the author studied it for the path algebroid $\mathcal{P}_{X}$ of a smooth variety $X$ and named it as formal path groupoid.
@@
For an $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra $(\mathcal{A},\Delta,\epsilon)$, the sheaf of premitive elements is determined by the following assignment (see [MS22])
@@
For a smooth ((((holomorphic)))) Lie algebroid [BMRT15] i.e. for a locally free sheaf of Lie-Rinehart algebras of finite rank over a smooth (complex) manifold, the universal enveloping algebroid is a cocomplete locally graded free $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra ((((of finite type)))) [MS22].
The tangent sheaf over the affine scheme $Spec(\mathbb{A}^{\mathbb{N}})$ ((((where $\mathbb{A}^{\mathbb{N}}:=\mathbb{K}[x_{i}]_{i\in\mathbb{N}})$ [Dri06, MS22] and the free Lie algebroid $\mathcal{P}_{X}$ over a smooth manifold $X$ [Kap07] are locally free sheaf of Lie-Rinehart algebras ((((of infinite type)))), the universal enveloping algebroid is a cocomplete locally graded free $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra ((((of infinite type)))) [MS22].
@@ 3. Lie algebroid Cohomology over $a$-spaces
Lie algebroids over algebraic spaces (shortly, $a$-spaces) is discussed in Section 2.1 by considering certain quasicoherent sheaves of Lie-Rinehart algebras over $a$-spaces $(X,\mathcal{O}_{X})$ (form an abelian category). The category of representations of a Lie algebroid has enough injectives (since the category of quasicoherent sheaves of left $\mathcal{O}_{X}$-modules is enough injectives [Har77]), useful for doing homological algebra.
The Chevalley-Eilenberg-de Rham cohomology (or, Lie algebroid cohomology) of a coherent Lie algebroid over a complex manifold or over a Noetherian separated scheme is expressed as derived functor $Ext$ [Pym13, Bru17]. In [Pym13], the result is stated for locally free Lie algebroid of finite rank over a complex manifold and in [Bru17], the result is proved for locally free Lie algebroid of finite rank over a Noetherian separated scheme. Here, we construct a more general cochain complexs of sheaves, which coincides with the cochain complexes of sheaves appear in the classical cases [Mac05, Pym13, BMRT15, Bru17]. For that we do not require the condition of coherent (or locally finitely presented) $\mathcal{O}_{X}$-module for a Lie algebroid. Note that for analytic and algebraic geometric setups, we do not need the base space $X$ to be non-singular. However, in that case, the tangent sheaf, which becomes a coherent Lie algebroid, does not retain the property of being a locally free Lie algebroid. For more details, see [Pym13, MS22]. Moreover, we present the result in a simplified form by using Čech cohomology. As an application of Lie algebroid cohomology we consider Chern classe of a module (locally free $\mathcal{O}_{X}$-module of finite rank) over the Lie algebroid.
@@ 3.1. Atiyah algebroid
For an (quasicoherent) $\mathcal{O}_{X}$-module $\mathcal{E}$, we form a Lie algebroid consists of the sheaf of first order differential operators on $\mathcal{E}$, defined by (generalizes the notion of Atiyah algebroids [Pym13, Bru17])
@@
(1) $0\rightarrow\mathscr{E}nd_{\mathcal{O}_{X}}(\mathcal{E})\hookrightarrow\mathcal{A}t(\mathcal{E})\overset{\mathfrak{a}}{\rightarrow}\mathcal{T}_{X}\rightarrow 0$
In particular, when $\mathcal{E}=\mathcal{O}_{X}$ for a non singular $a$-space $X$, we have $\mathcal{A}t(\mathcal{O}_{X})\cong\mathcal{O}_{X}\oplus\mathcal{T}_{X}$, equals to the sheaf differential operators of order $\leq 1$ with scalar symbol. The universal enveloping algebroid of the Atiyah algebroid $\mathcal{A}t(\mathcal{O}_{X})$ is the sheaf of twisted differential operators over $X$ [BP15, MS22].
A connection $\nabla:\mathcal{T}_{X}\rightarrow\mathcal{A}t(\mathcal{E})$ that satisfies $\mathfrak{a}\circ\nabla=Id_{\mathcal{T}_{X}}$ provides a splitting of the s.e.s. (1) as $\mathcal{O}_{X}$-modules. If its curvature is zero then the s.e.s. splits as Lie algebroids. More generally, for a Lie algebroid $(\mathcal{L},[\cdot,\cdot],\mathfrak{a})$ on an $a$-space $(X,\mathcal{O}_{X})$, a $\mathcal{L}$-connection on $\mathcal{E}$ is defined by an $\mathcal{O}_{X}$-linear map (see [Pym13, BP15, Bru17])
@@
satisfying the Leibniz rule $\nabla_{D}(f~{}s)=f~{}\nabla_{D}(s)+\mathfrak{a}(D)(f)~{}s$ for each local sections ((((for require compatibility, we choose $\sigma\circ\nabla=\mathfrak{a}$, i.e. $\sigma_{\Delta_{D}}=\mathfrak{a}(D)$ for $D\in\mathcal{L})$. It is equivalently described by a $\mathbb{K}_{X}$-linear map
@@
satisfying the Leibniz rule $\nabla(f~{}s)=f~{}\nabla s+\breve{\mathfrak{a}}(df)\otimes s$, where $\Omega_{\mathcal{L}}^{1}:=\mathscr{H}om_{\mathcal{O}_{X}}(\mathcal{L},\mathcal{O}_{X})$ for a Lie algebroid $\mathcal{L}$, $\Omega_{X}^{1}:=\Omega_{\mathcal{T}_{X}}^{1}$ and $\breve{\mathfrak{a}}:\Omega_{X}^{1}\rightarrow\Omega_{\mathcal{L}}^{1}$ is the dual of the anchor map.
@@
$(ii)$ For $\mathcal{L}=\mathcal{T}_{X}(-logY)$, if a connection exists on an $\mathcal{O}_{X}$-module $\mathcal{E}$, it is called a logarithmic connection on $\mathcal{E}$. It forms a meromorphic connection with poles along the divisor $Y$ [CJNMM96, Pym13].
@@
For a $\mathcal{O}$-module $\mathcal{F}$ on $X$, there exists an injective resolution $\mathcal{F}\overset{\sim}{\rightarrow}\mathcal{C}^{\bullet}(\mathcal{F})$ ((((a quasi-isomorphism) of $\mathcal{O}$-modules. Thus, for a complex of $\mathcal{O}$-modules $\mathcal{F}^{\bullet}$, we consider the associated bicomplex of sheaves $\mathcal{C}^{\bullet}(\mathcal{F}^{\bullet})=(\mathcal{C}^{p}(\mathcal{F}^{q}))$ $(p,q\in\mathbb{N}\cup\{0\})$ of injective resolutions. The original complex is embedded in the total complex $\mathcal{K}^{\bullet}=tot(\mathcal{C}^{\bullet}(\mathcal{F}^{\bullet}))$, and this embedding is a quasi-isomorphism. The cohomology of the associated complex $\mathcal{K}^{\bullet}(X)=tot(\mathcal{C}^{\bullet}(\mathcal{F}^{\bullet}))(X)$ of global sections is called the hypercohomology of $\mathcal{F}^{\bullet}$, and denoted by $\mathbb{H}^{\bullet}(X,\mathcal{F}^{\bullet})$ (see [Dan89]).
@@
The $k$-th hypercohomology is a functor $\mathbb{H}^{k}(X,-):Shv(X)^{\bullet}\rightarrow Vect$ for $k\in\mathbb{N}\cup\{0\}$ (see [Dan89, Ste15] for details) satisfying the following two conditions:
@@
We use sheaf-theoretic generalizations of the well-known derived functors $Ext$ and $Cotor$, by sheafifying the standard cochain complexes ((((see [Rin63], [KP11])))) accordingly and then considering the associated hypercohomologies.

### v7

@@ 2.2. Universal enveloping algebroid of a Lie algebroid
Let $(\mathcal{L},[\cdot,\cdot],\mathfrak{a})$ be a Lie algebroid over any of the special $a$-space $(X,\mathcal{O}_{X})$. For each open set $U$ of $X$, we find the universal enveloping algebra $\mathcal{U}(\mathcal{O}_{X}(U),\mathcal{L}(U))$ of the $(\mathbb{K},\mathcal{O}_{X}(U))$-Lie-Rinehart algebra $\mathcal{L}(U)$ or of the Lie-Rinehart pair $(\mathcal{O}_{X}(U),\mathcal{L}(U))$ (see [34, 16, 26]). The sheafification of the cannonical presheaf:
@@
is known as the universal enveloping algebroid of the Lie algebroid $\mathcal{L}$, and denoted by $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ (see [5, 32, 3, 35, 29] for details).
@@
The sheaf of logarithmic derivations and sheaf of logarithmic differential operators are denoted as $\mathcal{T}_{X}(-logY)$ and $\mathcal{D}_{X}(-logY)$ respectively for a principal divisor $Y$ in some complex manifold or smooth algebraic variety $X$ [29].
In the case of a free divisor $Y$ in $X$ (i.e. $\mathcal{T}_{X}(-logY)$ is locally free $\mathcal{O}_{X}$-module [4, 32, 29]) we have (sheafifying the local description given for the module of logarithmic derivations in [22]) the isomorphism
@@
In [18], the notion of path algebroid $\mathcal{P}_{X}$ of a smooth manifold or smooth algebraic variety $X$ is constructed as the free Lie algebroid generated by tangent sheaf $\mathcal{T}_{X}$. It forms a locally free sheaf of Lie-Rinehart algebras over $(X,\mathcal{O}_{X})$ of infinite rank. The universal enveloping algebroid $\mathscr{U}(\mathcal{O}_{X},\mathcal{P}_{X})=:\mathbb{D}_{X}$ is described as sheaf of non commutative differential operators on $X$.
@@ 2.3. $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebras
The notation $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebras [29] is considered as a sheaf theoretic analogue of the notion $R/{\mathbb{K}}$-bialgebras [26] or left bialgebroids [19]. As an example, we consider the universal enveloping algebroid $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ (and its dual) of a Lie algebroid $\mathcal{L}$ over $(X,\mathcal{O}_{X})$. Similar structure is appeared in [18, 7, 5, 2], in complex and algebraic geometry context.
@@
We recall the canonical $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra structures on $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ and its dual in the following.
@@
Observe that $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ is a cocommutative counital $\mathcal{O}_{X}$-coalgebra in the sense that there is a natural cocommutative coassociative comultiplication $\Delta:\mathscr{U}(\mathcal{O}_{X},\mathcal{L})\rightarrow\mathscr{U}(\mathcal{O}_{X},\mathcal{L})\otimes_{\mathcal{O}_{X}}\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$, and counit $\epsilon:\mathscr{U}(\mathcal{O}_{X},\mathcal{L})\rightarrow\mathcal{O}_{X}$ respectively, which are locally given by the following formulas [5]
@@
$\epsilon(\tilde{D})=\tilde{D}(1)$,
for a section $f$ of $\mathcal{O}_{X}$, for a section $D$ of $\mathcal{L}$ with $\bar{D}=\iota_{\mathcal{L}}(D)$, and for sections $D^{\prime},D^{\prime\prime},\tilde{D}$ of $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$. Consequently, $(\mathscr{U}(\mathcal{O}_{X},\mathcal{L}),\Delta,\epsilon)$ is a sheaf of $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebras.
@@
For each open set $U$ of $X$, we find the jet algebras $\mathcal{J}(\mathcal{O}_{X}(U),\mathcal{L}(U))$ of the $(\mathbb{K},\mathcal{O}_{X}(U))$-Lie-Rinehart algebra $\mathcal{L}(U)$ (see [7, 19, 2]). The jet algebroid $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$ is the sheafification of the canonical presheaf:
@@
The product on $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ descends to a sheaf homomorphism $m:\mathscr{U}(\mathcal{O}_{X},\mathcal{L})\otimes_{\mathbb{K}_{X}}\mathscr{U}(\mathcal{O}_{X},\mathcal{L})\rightarrow\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$. We can therefore dualize the product to obtain a coproduct $\Delta^{*}:\mathscr{J}(\mathcal{O}_{X},\mathcal{L})\rightarrow\mathscr{J}(\mathcal{O}_{X},\mathcal{L})\otimes_{\mathcal{O}_{X}}\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$ locally defined as
@@
The jet algebroid is a formal groupoid that serves as the formal exponentiation of the Lie algebroid [5]. In [18], the author studied this structure for the path algebroid $\mathcal{P}_{X}$ of a smooth variety $X$, referring to it as the formal path groupoid.
@@
For an $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra $(\mathcal{A},\Delta,\epsilon)$, the sheaf of premitive elements is determined by the following assignment (see [29])
@@
For a smooth ((((holomorphic)))) Lie algebroid [1] i.e. for a locally free sheaf of Lie-Rinehart algebras of finite rank over a smooth (complex) manifold, the universal enveloping algebroid is a cocomplete locally graded free $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra ((((of finite type)))) [29].
The tangent sheaf over the affine scheme $Spec(\mathbb{A}^{\mathbb{N}})$ ((((where $\mathbb{A}^{\mathbb{N}}:=\mathbb{K}[x_{i}]_{i\in\mathbb{N}})$ [9, 29] and the free Lie algebroid $\mathcal{P}_{X}$ over a smooth manifold $X$ [18] are locally free sheaf of Lie-Rinehart algebras ((((of infinite rank)))). Thus, their universal enveloping algebroids are cocomplete locally graded free $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebras ((((of infinite type)))) [29].
@@ 3. Lie algebroid Cohomology over $a$-spaces
Lie algebroids over algebraic spaces (shortly, $a$-spaces) are discussed in Section 2.1 by considering certain quasicoherent sheaves of Lie-Rinehart algebras over $a$-spaces $(X,\mathcal{O}_{X})$ (form an abelian category). The category of representations of a Lie algebroid has enough injectives (since the category of quasicoherent sheaves of left $\mathcal{O}_{X}$-modules is enough injectives [14]), useful for doing homological algebra.
The Chevalley-Eilenberg-de Rham cohomology (or, Lie algebroid cohomology) of a coherent Lie algebroid over a complex manifold or over a Noetherian separated scheme is expressed as derived functor $Ext$ [32, 3]. In [32], the result is stated for locally free Lie algebroid of finite rank over a complex manifold and in [3], the result is proved for locally free Lie algebroid of finite rank over a Noetherian separated scheme. Here, we construct a more general cochain complex of sheaves, which coincides with the cochain complexes of sheaves used in the classical contexts [21, 32, 1, 3]. For that we do not require the coherent (or locally finitely presented) condition on the underlying $\mathcal{O}_{X}$-module of a Lie algebroid. Notably, for analytic and algebraic geometric setups, the base space $X$ does not need to be non-singular. However, in that case, the tangent sheaf, which becomes a coherent Lie algebroid, does not retain the property of being a locally free Lie algebroid. For more details, see [32, 29]. Moreover, we present the result in a simplified form by using Čech cohomology. As an application of Lie algebroid cohomology we consider Chern classe of a module (locally free $\mathcal{O}_{X}$-module of finite rank) over the Lie algebroid.
@@ 3.1. Atiyah algebroid
For an (quasicoherent) $\mathcal{O}_{X}$-module $\mathcal{E}$, we form a Lie algebroid consisting of the sheaf of differential operators on $\mathcal{E}$ of order $\leq 1$ [18, 32, 3], defined by
@@
[…]
In particular, when $\mathcal{E}=\mathcal{O}_{X}$ for a non singular $a$-space $X$, we have $\mathcal{A}t(\mathcal{O}_{X})\cong\mathcal{O}_{X}\oplus\mathcal{T}_{X}$, equals to the sheaf differential operators of order $\leq 1$ with scalar symbol. The universal enveloping algebroid of the Atiyah algebroid $\mathcal{A}t(\mathcal{O}_{X})$ is the sheaf of twisted differential operators over $X$ [2].
A connection $\nabla:\mathcal{T}_{X}\rightarrow\mathcal{A}t(\mathcal{E})$ that satisfies $\sigma\circ\nabla=Id_{\mathcal{T}_{X}}$ provides a splitting of the s.e.s. (1) as $\mathcal{O}_{X}$-modules. If its curvature is zero then the s.e.s. splits as Lie algebroids. More generally, for a Lie algebroid $(\mathcal{L},[\cdot,\cdot],\mathfrak{a})$ on an $a$-space $(X,\mathcal{O}_{X})$, a $\mathcal{L}$-connection on $\mathcal{E}$ is defined by an $\mathcal{O}_{X}$-linear map (see [32, 2, 3])
@@
satisfying the Leibniz rule $\nabla_{D}(f~{}s)=f~{}\nabla_{D}(s)+\mathfrak{a}(D)(f)~{}s$ for sections $f\in\mathcal{O}_{X}$, $D\in\mathcal{L}$ and $s\in\mathcal{E}$ ((((to ensure compatibility, we choose $\sigma\circ\nabla=\mathfrak{a}$, i.e. $\sigma_{\nabla_{D}}=\mathfrak{a}(D)$ for $D\in\mathcal{L})$. Equivalently, it is described by a $\mathbb{K}_{X}$-linear map
@@
satisfying the Leibniz rule $\nabla(f~{}s)=f~{}\nabla s+\breve{\mathfrak{a}}(df)\otimes s$, where $\Omega_{\mathcal{L}}^{1}:=\mathscr{H}om_{\mathcal{O}_{X}}(\mathcal{L},\mathcal{O}_{X})$ for a Lie algebroid $\mathcal{L}$ and $\Omega_{X}^{1}:=\Omega_{\mathcal{T}_{X}}^{1}$ together with $\breve{\mathfrak{a}}:\Omega_{X}^{1}\rightarrow\Omega_{\mathcal{L}}^{1}$ is the dual of the anchor map.
@@
$(ii)$ For $\mathcal{L}=\mathcal{T}_{X}(-logY)$, if a connection exists on an $\mathcal{O}_{X}$-module $\mathcal{E}$, it is called a logarithmic connection on $\mathcal{E}$. It forms a meromorphic connection with poles along the divisor $Y$ [4, 32].
@@
For a $\mathcal{O}$-module $\mathcal{F}$ on $X$, there exists an injective resolution $\mathcal{F}\overset{\sim}{\rightarrow}\mathcal{C}^{\bullet}(\mathcal{F})$ ((((a quasi-isomorphism) of $\mathcal{O}$-modules, known as the flabby Godement resolution of $\mathcal{F}$. Thus, for a complex of $\mathcal{O}$-modules $\mathcal{F}^{\bullet}$, we consider the associated bicomplex of sheaves $\mathcal{C}^{\bullet}(\mathcal{F}^{\bullet})=(\mathcal{C}^{p}(\mathcal{F}^{q}))$ $(p,q\in\mathbb{N}\cup\{0\})$ of injective resolutions. The original complex is embedded in the total complex $\mathcal{K}^{\bullet}=tot(\mathcal{C}^{\bullet}(\mathcal{F}^{\bullet}))$, and this embedding is a quasi-isomorphism. The cohomology of the associated complex $\mathcal{K}^{\bullet}(X)=tot(\mathcal{C}^{\bullet}(\mathcal{F}^{\bullet}))(X)$ of global sections is called the hypercohomology of $\mathcal{F}^{\bullet}$, and denoted by $\mathbb{H}^{\bullet}(X,\mathcal{F}^{\bullet})$ (see [8]).
@@
The $k$-th hypercohomology is a functor $\mathbb{H}^{k}(X,-):Shv(X)^{\bullet}\rightarrow Vect$ for $k\in\mathbb{N}\cup\{0\}$ (see [8, 36] for details) satisfying the following two conditions:
@@
We use sheaf-theoretic generalizations of the well-known derived functors $Ext$ and $Cotor$, by sheafifying the standard cochain complexes ((((see [34], [19])))) and subsequently considering the associated hypercohomologies.